# Visual Affordances: Enabling Robots to Understand Object Functionality

### v1

@@ II-F End-effector pose estimation and synthesis
Methods for end-effector pose estimation and synthesis focus on a generic grasping interaction with the objects in the image, without taking into account the task that the robot performs and the affordances that the object supports. This fact can result in estimating or synthesising poses that are not aligned with the task, leading to interactions with the object that do not assist the human.
II-G Future directions: AI agents and human-in-the-loop
Extension of our perception-robotic framework to multi-modal inputs, and especially the use of language, can lead to its re-design as an agentic system [108] [^4] https://www.anthropic.com/engineering/building-effective-agents
https://blogs.nvidia.com/blog/what-is-agentic-ai/
. The framework requires steps that are commonly refer to as understanding (perception part), reasoning (relating affordances, objects, and physical properties while being conditioned to the task to accomplish), planning (actuation part of the robot to accomplish the task), and recovering from errors when referring to AI agents. One way to recover from errors is to feed the model with previous predictions to correct the mistakes [109]. Learning to predict visual affordances for hand-object interactions can benefit from human demonstrations of the actions to perform, in the same way humans prompt models with examples showing how to solve tasks [109, 110]. Our framework can be extended to include the feedback from a person at different stages (human-in-the-loop) [111, 112]. For example, by correcting the prediction mistakes of single models or also by injecting task specific knowledge in the process, overriding a specific model. Human behaviours can be collected to update the models and improve the interaction with objects.
@@ II-G Future directions: AI agents and human-in-the-loop
KEYS – # Images: number of images, OBJ: number of object categories, AFF: number of affordance categories, Tran.: transparency, 3PV: third person view, HOc: hand-occlusion, OBJD: task driven object detection; AFFC: affordance classification; AFFG: affordance grounding; HOIS: hand-object pose estimation and interaction synthesis; GDET: grasping detection; AFFDS: affordance detection and segmentation; : considered, : not considered, : partially considered.
III Datasets: review and limitations
In this section, we compare the characteristics of image-based datasets for visual affordance prediction and discuss their similarities and limitations (see Table VII), contrary to previous surveys [26, 28]. Our comparison considers
  • the type of environment (indoor or outdoor);
  • the camera viewpoint (third person or first person);
  • the objects of interest (quantity, diversity depending on the group such as tools or containers, physical properties such as transparency);
  • the type of images (real, simulated, mixed-reality);
  • the presence of occlusions, caused by other objects, clutter, or hand manipulating the object; and
  • the annotations of affordances (quantity, accuracy, procedure, and expertise of the annotators).
These datasets are usually split into two non-overlapping sets: one to train learning-based models (training set) and another to evaluate the performance of the models (testing set). Dataset choices, biases in the images, and ambiguities or inaccuracies in the annotations affect and are transferred to the models during the training phase.
Annotations of affordances. Previous works proposing datasets either reuse images already available for other tasks such as object detection or image classification [11, 12, 14, 16, 17, 20, 98], or collect new images specifically for visual affordance prediction [9, 13, 33]. Target affordances in most of these datasets are manually labelled. For example, affordance segmentation requires the annotations to label the pixels of the object regions with an affordance class (fine-grained annotation) [9, 12, 16, 33, 114, 115, 116]. However, this procedure is time-consuming and subject to errors (see Fig. 6).
Figure 6: Examples of manual annotation on real data from UMD [9] and IIT-AFF [12]. First row: RGB samples. Second row: manual annotation of affordance prediction (segmentation). Mistakes highlighted with orange rectangles: not every object in the image is annotated (1st column), missing regions (2nd and 4th column), or the annotation of the graspable handle is overlapped with the hand occluding the object (3rd column).
@@ III Datasets: review and limitations
Only some objects are annotated, whereas other objects are ignored because of clutter in the scene or their size in the image is small due to the distance of the object from the camera [9, 12]. Annotated segmentation masks can be incomplete (presence of holes) or not accurate close to the object boundaries. To reduce the annotation effort, a weakly labelling procedure requires annotators to only label points of interaction and then applies a Gaussian blur operation on the image to expand the point annotation [11, 98]. This procedure was used to annotate two datasets for affordance grounding, OPRA [98] and AGD20K [11]. The blurring operation however may cause the probability of interaction to be non-zero also outside the object boundaries. Because of ambiguities in the boundaries of visual affordances and fine-grained annotations, size of the datasets are often limited to few tens of thousands. To scale the size of the datasets, simulators can generate a large number of synthetic or mixed-reality images with automatic annotations while varying the illumination conditions and object models [20, 106, 114, 115]. The task-dependent effort requires the design of the simulated environments, the placement of the objects CAD model in the environment, and the annotation of the affordances (e.g. the manual labelling of the mesh or point cloud with the affordance category) [13, 114, 115, 120]. The affordance annotations are then easily rendered on the camera frame in the simulator. For example, parts of the object mesh are associated with an affordance category and segmentation masks are obtained by ray-tracing the annotated mesh into the simulated camera frame [20, 114, 115]. For grasping detection, a robotic hand grasping the object can be simulated and the simulating tool can save the image of the object, the coordinates of the grasping attempts, and the bounding boxes as an annotation of the visual affordance [106]. However, images generated with a simulator can still differ from images captured with a real camera (sim-to-real gap), hindering the generalisation of trained models to real images, when deployed on a real robot. As an alternative to the use of simulators, an (semi-)automatic annotation procedure can use off-the-shelf models. For example, HANDAL [13] was annotated by using BundleSDF [121], a method that estimates the 6D pose of the objects in each frame of a video and reconstructs their CAD models, followed by manually labelling the handle of the CAD models with the affordance graspable, and projecting the annotated CAD model in the camera frame to obtain the annotation mask. When collecting HO3Pairs [5] to perform the synthesis of visual affordances (hand-object interactions) from egocentric images, the problem was to obtain an image containing only the object from a dataset of hand occluded objects (HOI4D [122]). In this case, annotators segmented the hand and used an image in-painter [123] to erase the hand holding the objects and reconstruct the occluded part of the object. Although the in-painting allows to obtain the image without the hand, the reconstruction causes the image quality to degrade with the presence of blurred areas, potentially affecting the performance of the trained method. The dataset RIO [113] was annotated by providing a description of the objects in the scene and an image as a prompt to an LLM, such as ChatGPT [124], that generated the description of the potential interaction between a human and the object. Using ChatGPT to automatically generate annotations requires annotators to setup the appropriate prompt and check for potential mistakes in the labels. Overall, using off-the-shelf methods can speed-up the labelling procedure, potentially scaling the size of annotated datasets [125].
Figure 7: Images from visual affordance prediction datasets [9, 12, 13, 16, 17, 33, 114, 115, 116]. Datasets have different background conditions, from single object on a plain coloured background to clutter or hand-occluded objects. Images are resized at the same height keeping the aspect ratio for visualization purpose.
Camera viewpoint: third person view and egocentric view. The majority of datasets for affordance prediction focuses on images from third person perspective [9, 11, 12, 13, 33, 98, 103, 106, 114, 116, 117]. The camera is not mounted on a robot or worn by a person, but it is in a fixed position, capturing objects from a constant distance and in a static scene (see Fig. 7). In some cases [103, 106, 117], the camera is placed in a top-down view to observe the area around a robotic arm. These conditions can limit the models trained on these dataset to generalise to other scenarios, e.g. when the view is not top-down or when there is motion blur in dynamic scenes. The first person perspective (egocentric view) includes additional challenges, such as self-occlusions due to the presence of parts of the human/robot in the collected frames and image blur due to the camera movement [5, 17, 52, 119]. For example, arms are observed from the bottom of an image resulting in objects highly occluded by the hands (e.g. images in FPHA-AFF [17]), or images are affected by blur while people are using kitchen tools and interacting with ingredients in a kitchen environment (e.g. EPIC-Kitchens [5, 119]). Because of these challenges, models trained on egocentric-view datasets might not generalise on images from third person perspective and vice versa.
Occlusions. Most of the datasets [5, 9, 103, 106, 115, 116, 117, 126] focus on one unoccluded object placed on a flat surface (e.g. tabletop or floor), and the fixed setup enabled the collection of images with a diverse and large number of object categories and objects instances. For example, UMD [9] and Multi-View [116] collected more than 15 object categories and annotated more than 5 affordance classes, while controlling the environmental conditions for all the images: same illumination and background, and objects placed on a rotating table. However, this simple and controlled setup limits the generalisation of models trained on these datasets to environments with different illumination and backgrounds, or where multiple objects are present in the scene. Only some of the datasets [11, 12, 13, 14, 16, 20] have images with occlusions caused by clutter in the scenes or human hands holding the objects (hand-occlusions). When objects are occluded, only some of their regions are visible, increasing the difficulty of learning affordance prediction because models can predict affordances only from partial information. Hand-occlusion is a main challenge in human-robot collaborations, as erroneous or inaccurate affordance predictions lead to unintended interactions with the object, potentially causing harm to the person (human safety[13, 20, 127].
35353535404040400102030405060051015Number of object categoriesNumber of affordance categoriesAGD20KYCB-AffordanceMulti-viewTRANS-AFFCHOC-AFFPieropan et al.Kjellström et al.Sun et al.Zheng et al.IIT-AFFCAD120-AFFFPHA-AFFGraspSegHANDALUMD-SynthUMDAFF-Synth
@@
Figure 8: Visualisation of datasets size based on number of images, number of affordance categories, and number of object categories. Note that we considered only datasets with available information. KEY:  Affordance classification,  Affordance detection and segmentation,  Affordance grounding,  Hand-object pose estimation.
@@
Objects of interest. In previous works [9, 11, 12, 13, 14, 16, 33, 103, 116, 117], the objects most suitable to accomplish a task are considered as objects of interest, and are associated with the corresponding affordances. Fig. 8 shows the size of the datasets based on the number of object and affordance categories. The majority of these datasets [3, 9, 12, 13, 16, 17, 54, 58] have fewer than 20 object categories and 10 affordance categories. Reported datasets focus on the affordances of tools and containers. Tools are usually opaque and rigid objects used in a kitchen environment (e.g. pan, fork, knife, turner) or for carpentry (e.g. hammer, shovel, saw) and consist of a graspable handle [9, 12, 13]. Compared to tools, perceiving the affordance of containers (e.g. box, cup, glasses) is more challenging, since their properties can change during a manipulation (e.g. the mass due to emptying or filling, or their appearance in case of the transparent material filled with opaque content) [18, 20, 120]. Even if a lot of containers we use in everyday life are transparent, this property is considered only in a few datasets [12, 19, 20, 33].
@@
Diversifying the object categories, degrees of occlusions and object poses in datasets is fundamental to tackle the generalisation problem. The generalisation to diverse conditions is particularly relevant in assistive applications, where the environment is not necessarily controlled and the robot’s objective is to help a human in a safe way.
III-A Future direction: scaling visual affordance datasets

### v2

@@ III-D Hand pose estimation and synthesis
Methods for multi-finger pose estimation and synthesis focus on a generic grasping interactions, without taking into account the task that the agent performs and the affordances that the object supports. This fact can result in estimating wrong poses not aligned with the task.
IV Datasets: review and limitations
In this section, we compare the characteristics of image-based datasets for visual affordance prediction and discuss their similarities and limitations (see Table VII), contrary to previous surveys [19, 21]. Our comparison considers elements such as: the type of environment (indoor or outdoor); the camera viewpoint (third person or first person); the objects of interest (quantity, diversity depending on the group such as tools or containers, physical properties such as transparency); the type of images (real, simulated, mixed-reality); the presence of occlusions, due to clutter, or hand manipulating the object; and the annotations of affordances (quantity, accuracy, procedure, and expertise of the annotators). These datasets are usually split into two non-overlapping sets: one to train models (training set) and another to evaluate their performance (testing set). During training, biases in the images, ambiguities or inaccuracies in the annotations are transferred to the models.
@@ IV Datasets: review and limitations
KEYS – # Images: number of images, OBJ: number of object categories, AFF: number of affordance categories, Tran.: transparency, 3PV: third person view, HOc: hand-occlusion, OBJD: task driven object detection; AFFC: affordance classification; AFFG: affordance grounding; HOIS: hand-object pose estimation and interaction synthesis; GDET: grasping detection; AFFDS: affordance detection and segmentation; : considered, : not considered, : partly considered.
@@
Annotations of affordances. Previous works proposing datasets either sampled images already available for other tasks such as object detection or image classification [6, 8, 10, 13, 56, 59, 93], or collected new images [7, 11, 26]. Target affordances in most of these datasets are manually labelled. For example, affordance segmentation requires to label the pixels of the object regions with an affordance category (fine-grained annotation) [6, 10, 11, 26, 108, 109, 110]. However, this procedure is time-consuming and subject to errors, such as missing annotations of objects  [11, 6], incomplete annotation (presence of holes) or over the object boundaries, due to clutter or small visible regions. To reduce the annotation effort, a weakly labelling procedure requires annotators to only label points of interaction and then to apply a Gaussian filter on the image to expand the point annotation [13, 93]. This procedure was used to annotate two datasets for affordance grounding, OPRA [93] and AGD20K [13]. The filtering operation however may cause the affordance map to be non-zero also outside the object boundaries. Because of ambiguities in the boundaries of visual affordances and fine-grained annotations, datasets size are often limited to few tens of thousands images. Simulators can generate a large number of synthetic or mixed-reality images with automatic annotations while varying the illumination conditions and object models [59, 105, 108, 109]. In this case, the annotation effort consists in the design of the simulated environments, the placement of the object models, and the manual labelling of the mesh with the affordance category [7, 59, 108, 109]. Segmentation masks are obtained by ray-tracing the annotated regions on the object mesh into the simulated camera frame [59, 108, 109]. A robotic hand grasping the object can be simulated to save the image of the object, the coordinates of the grasping attempts, and the oriented rectangles [105]. However, images generated with a simulator can differ from images captured with a real camera (sim-to-real gap), hindering the generalisation of trained models to real images. An alternative to simulators, is a (semi-)automatic annotation procedure using off-the-shelf models [5, 7, 107, 111]. For example, HANDAL [7] was annotated by using BundleSDF [115] to estimate the 6D pose of the objects in each frame of a video and to reconstruct their CAD models. Then, the handle of the CAD models were annotated with the affordance graspable, and projected in the camera frame to obtain the annotation mask. EPIC-AFF [111] annotation procedure associated the action narrations from EPIC-100 [114] with the hand-object interaction points from VISOR [116], and projected these points in 3D using a depth estimation model [117]; finally, COLMAP [118] estimated the camera poses and projected interaction points in the same environment point cloud, causing the affordance regions to cover the object, the background, and the arms of the person when projected in the camera frames. When collecting HO3Pairs [5] to perform the synthesis of visual affordances, annotators segmented the hand and used an image in-painter [119] to erase the hand holding the objects reconstructing the occluded part of the object. Although this procedure allows to obtain the image showing the unoccluded object, the reconstruction causes the image quality to degrade with the presence of blurred areas, affecting the performance of trained methods. UMD VL and IIT-AFF VL [85] complemented the existing UMD and IIT-AFF annotations with two language description of the task: explicit instructions included object category, action verb, and positional relationship (e.g., “Hand me the [object] on the right to [action]”), and implicit instructions omitting specific object category references (e.g., “Hand me something to [action]”). CLIP [35] predicted the ”[object]” category from the crops of detected objects, while GPT-2 [120] generated the ”[action]” from the template of the instruction to fill. Overall, using off-the-shelf methods can speed-up the labelling procedure, potentially scaling the size of annotated datasets [121], requiring annotators to setup the annotation pipeline and check for potential mistakes in generated labels.
@@
Figure 3: Visualisation of datasets size based on number of images, number of affordance categories, and number of object categories. KEY:  Affordance classification,  Affordance detection and segmentation,  Affordance grounding,  Hand-object pose estimation.
Camera viewpoint: 3rd and 1st person view. The majority of datasets for visual affordance focuses on third person view [6, 7, 11, 13, 26, 93, 102, 105, 108, 110, 112]. The camera has a fixed pose and is not mounted on the agent, capturing objects from a constant distance and in a static scene. In some cases [102, 105, 112], the camera is placed in a top-down view to observe the area close to the agent. These conditions can limit the generalisation of trained models to other scenarios, e.g. different camera view. The first person perspective (egocentric view) includes additional challenges, such as self-occlusions due to the presence of parts of the agent in the collected frames and blur due to the camera movement [5, 44, 56, 114]. For example, arms are observed from the bottom of an image resulting in objects highly occluded by the hands (e.g. FPHA-AFF [56]), or images are affected by blur while people interact with ingredients in a kitchen environment (e.g. EPIC-Kitchens [5, 111, 114]). Because of these challenges, models trained on egocentric-view datasets might not generalise to third person perspective and vice versa.
Occlusions. Most of the datasets [5, 11, 102, 105, 109, 110, 112, 122] focus on one unoccluded object placed on a flat surface (e.g. tabletop or floor) with a fixed setup varying object categories and objects instances. For example, UMD [11] and Multi-View [110] collected more than 15 object categories and annotated more than 5 affordance classes, while controlling the environmental conditions: objects are placed on a rotating table, with the same illumination and background. However, this simple and controlled setup limits the generalisation of models to environments with different illumination and backgrounds, or where multiple objects are present in the scene. Only some of the datasets [6, 7, 8, 10, 13, 59] contain occlusions caused by clutter in the scenes or human hands holding the objects (hand-occlusions). When objects are occluded, only some of their regions are visible, increasing the difficulty in perceiving the affordances. Hand-occlusion is a main challenge in human-robot collaborations, as erroneous or inaccurate affordance predictions lead to unintended interactions with the object, potentially causing harm to the person (human safety[30, 59].
@@
Objects of interest. In previous works [6, 7, 8, 10, 11, 13, 26, 102, 110, 112], the objects most suitable to accomplish a task were considered as objects of interest, and were annotated with the corresponding affordances. The majority of these datasets [3, 6, 7, 10, 11, 46, 50, 56] has fewer than 20 object categories and 10 affordance categories (see Fig. 3), and focuses on the affordances of tools and containers. Tools are usually opaque and rigid, are used in a kitchen environment (e.g., pan, fork, turner) or for carpentry (e.g., hammer, shovel, saw), and consist of a graspable handle [6, 7, 11]. Compared to tools, perceiving the affordance of containers (e.g. box, cup, glasses) is more challenging, since their properties can change during a manipulation (e.g. the appearance in case of transparent material filled with opaque content) [59, 123, 124]. Even if a lot of containers we use in everyday life are transparent, this property is considered only in a few datasets [6, 26, 59, 97].
@@
Diversifying the object categories, degrees of occlusions and object poses in datasets is fundamental to tackle the generalisation problem. The generalisation to diverse conditions is relevant in human-robot collaboration and assistive applications, where the environment is not necessarily controlled.
V Affordance prediction reproducibility
We discuss the evaluation of affordance prediction, focusing on the reproducibility [^2] Principle of obtaining the same results given the same conditions (i.e. data, training and testing setups, and trained model) [125]. issues of current benchmarks, resulting in unfair and inconsistent comparisons. Reproducibility allows fair comparisons across methods and helps build upon previous works while understanding their limitations. We then highlight Open Science practices for fair benchmarking.
V-A Reproducibility challenges
  1. 1.
    data availability for benchmarking (RC1);
  2. 2.
    availability of a method’s implementation (RC2);
  3. 3.
    availability of trained models (RC3);
  4. 4.
    details of experimental setups (RC4); and
  5. 5.
    details of performance measures for evaluation (RC5).
In visual affordance prediction, no dataset is collected exclusively for benchmarking methods under specific conditions, such as illumination, clutter, or hand-occlusion (RC1). The majority of previous works [7, 11, 54, 59, 110] trained methods on a dataset training split and compared their performance on the testing split of one or more datasets. Cross-dataset evaluations are mostly avoided due to partial overlapping of affordance classes or of object categories, across the selected datasets [4]. For example, datasets such as UMD [11], IIT-AFF [6], and Multi-View [122] share some of the object and affordance classes, but labelled with different conventions, making the comparison of models trained on different dataset difficult. As a consequence, researchers train multiple version of the same model, adapted to the classes of a specific dataset. Additional documentation, such as metadata, help researchers train or evaluate methods only on common categories by re-ordering them. Moreover, relying only on a single benchmark can lead to limited and not generalisable considerations on model rankings. For example, images in UMD and Multi-View are collected in a laboratory environment with static conditions, such as a fixed camera oriented towards a table where an object is placed (camera-object distance is almost always the same) [11, 110]. However, in real scenarios the camera might be closer or farther from objects compared to the training setting, hence the performance on the benchmark might not reflect the performance on a real use case.
The lack of publicly available implementation of methods (RC2) [68, 72, 77, 78], the lack of publicly available trained models (RC3) [6, 54, 68, 72, 75, 77, 78], and the lack of details of experimental setups (RC4) [12, 54, 68, 72, 77, 78] can challenge researchers in reproducing previous works for comparative evaluations. The release of the model trained weights, of the method and inference pipeline implementation, is a crucial aspect for reproducibility, especially for deep-learning based models, allowing other researchers to test models on their own data without re-training. The availability of model implementation and weights is important when researchers need a comparison, as re-training the model can be too time- and resource-consuming. In case a new dataset is proposed and a previous method needs re-training, only the method implementation is sufficient. The re-implementation of methods and setup is time-consuming and prone to errors, and not always leads to the expected outcome (i.e. results are not replicable or findings are not reproducible). To avoid this issue and to save time, researchers report the results from previous works [68, 72, 77, 78], resulting in unfair comparisons if the experimental conditions are not the same, and in misleading findings and conclusions.
TABLE VIII: Comparison of training/testing setups used by different methods for affordance detection and segmentation on the UMD dataset [11]. Due to the setup inconsistencies, direct comparison among models performance is unfair.
Training setup Resolution Data augmentation Image resize
FLIP SCAL ROT JIT Train. Test.
AffordanceNet [12] $1000\times 600$ $\circ$ $\circ$ $\circ$ $\circ$ UNK UNK
CNN [54] $320\times 240$ $\circ$ $\circ$ $\circ$ $\circ$ $C$ SLW
DRNAtt [72] $320\times 240$ $\circ$ $\circ$ $\circ$ $\circ$ $C$ UNK
RANet [78] $224\times 224$ $\circ$ $\circ$ $\circ$ $\circ$ $C$ UNK
GSE [68] $400\times 400$ $\bullet$ $\bullet$ $\circ$ $\circ$ crop UNK
BPN [77] $1000\times 600$ $\bullet$ $\bullet$ $\bullet$ $\bullet$ UNK UNK
KEYS – $\bullet$: considered, $\circ$: not considered, FLIP: flipping, SCALE: scaling, ROT: rotating, JIT: colour jittering, Train.: training set, Test.: testing set, UNK: unknown, cc: centre-crop, SLW: sliding window
Using the same experimental setup to train and test affordance models allows a fair comparison enabling the validation of the technical contributions proposed by a novel work. When releasing the training and testing code is not possible, reporting all details to reproduce a setup becomes fundamental, enabling other researchers to re-implement the setup and correctly compare their solution. The experimental setup details include training hyper-parameter values, chosen data splits, image pre-processing (normalisation and cropping procedures), and post-processing. The lack of details of the experimental setup causes methods for affordance detection and segmentation to be often not reproducible [12, 54, 68, 77, 72, 78]. For example, AffordanceNet and BPN do not include image resize during training and testing phases [12, 77], whereas DRNAtt, RANet, and GSE do not include these details during the testing phase. Other details often omitted are the parameters of the optimizers used during training [72, 84, 88, 89]. Apicella et al.’s work [4] showed that the lack of details in the experimental setup led to unfair and inconsistent comparisons.
Previous works evaluated the performance of different methods using scores or metrics to quantify the discrepancy between predictions and annotations (more details in Supp. Mat.). Describing a performance measure help other researchers understand if the experiment validates their claim or if a different measure should be chosen. Providing the mathematical formulation of the scores helps disambiguate similar meaning but different implementations, especially when a public evaluation toolkit is not used or referred to. For example, mean IoU can be the average of all the IoUs between prediction and annotation, or the IoU considering the full set of predictions and annotations. Previous works evaluated a few methods with different performance measures or datasets, making comparison and ranking not possible. For example, the performance of AdaptiveNet [126] and STRAP [80] was compared on CAD120-AFF using IoU, instead of UMD using $F^{w}_{\beta}$ as most of available methods.
Affordance detection and segmentation methods are difficult to reproduce due to missing implementation and lack of setups details [12, 54, 68, 72, 77, 78]. We report the training and testing setups of affordance detection and segmentation methods on the UMD dataset in Table VIII. Despite being trained and tested on the same dataset, models’ performance is not directly comparable due to inconsistencies in the setups such as the image resize procedure (image cropping or input resolution) and augmentation procedure during training.
Inconsistencies can also be present in previous methods adapted into a baseline to compare with. For example, due to the missing annotation of the object pose in the training/testing dataset, AffordanceDiffusion [5] is compared with the coarse hand prediction of GanHand [8]. However, since a part of the architecture and of the training procedure is missing, the result is only a proxy to the (unknown) performance of GanHand.
TABLE IX: Affordance Sheet, inspired by Model Cards [23], to favour transparency and reproducibility of works for visual affordance predictions conditioned on robotic tasks. Example filled with ACANet [59] details.
ACANet
Affordance task OBJL FUNC FUNS EPE EIS
Datasets
Name: CHOC-AFF
Record link*: https://doi.org/10.5281/zenodo.5085800
Licence: CC BY 4.0
Proposed method
(RC2, RC3)
Record link*: https://doi.org/10.5281/zenodo.8364196
Code link: https://github.com/apicis/aff-seg/
Model card:
Licence: CC BY-NC-SA 4.0
Experimental setup
Data splits:
Set Images
Training 89,856
Validation 17,280
Testing 1 13,824
Testing 2 17,280
Hyperparameters:
Name Value
batch size 2
learning rate 0.001
schedule 0.5x
patience 3
optmizer SGD
momentum 0.9
weight decay 0.0001
resize [1, 1.5]
flip 0.5
Resize procedure: center crop $480\times 480$
Performance measures
Description: Per-class Jaccard index measures the overlap between predicted and annotated segmentation masks, and quantifies how much they are similar in size
Formulation: $\frac{\sum_{n=1}^{N}\sum_{\boldsymbol{y}\in I_{n}}TP^{\boldsymbol{y}}_{n}}{\sum_{n=1}^{N}\sum_{\boldsymbol{y}\in I_{n}}TP^{\boldsymbol{y}}_{n}+FP^{\boldsymbol{y}}_{n}+FN^{\boldsymbol{y}}_{n}}\>$
Limitations: Jaccard Index does not consider the masks shape
Robot validation Robot model: -
End-effector: -
Experiment: -
Legend: OBJL: object localisation; FUNC: functional classification; FUNS: functional segmentation; EPE: hand pose estimation; EIS: hand interaction synthesis; RC: reproducibility challenge. Notes: *data and weights of the trained model are recommended to be placed in a repository that favours long-term persistence and accessibility.
The redefinition of the visual affordance problem (see Sec. III) can also result in experimental validations ignoring datasets and benchmarks of partially overlapping formulations. For example, works on affordance grounding [13, 84, 88, 90] do not compare the performance of proposed methods with that of affordance segmentation methods [54, 68, 72, 78], even if the problem formulation is similar [13, 54]. Methods for affordance segmentation output a binary mask for each action in a predefined set of classes, whereas methods for affordance grounding output a confidence map describing where an action known a priori can take place in the image. Despite these differences, comparing methods for both affordance grounding and affordance detection and segmentation can explain if using action as input (affordance grounding) to a model provides any advantage.
V-B In support of reproducibility: Affordance Sheets
To promote reproducibility in affordance prediction, we propose the Affordance Sheet, an organised collection of good practices favouring fair comparisons and the development of new solutions (see Table XIII). Model cards [23] were previously introduced to improve the methods transparency and to raise awareness about limitations, by describing the method, the experimental setup, and the applications or conditions leading to underperformance. Our Affordance Sheet integrates Model Cards complementing the released information.
The first section identifies which problems the affordance model tackles, helping researchers understand what are the competing methods and assess their performance of solutions under the same inputs and conditions. When proposing a new problem partially overlapping with another one, previous models can be used or adapted to validate the method. For example, selecting the channel of an affordance segmentation output based on the action considered by the affordance grounding method enables the comparison between methods for affordance segmentation and methods for affordance grounding. To compare the grounding and segmentation outputs, the grounding confidence map can be converted to a binary mask via thresholding; alternatively, the segmentation map can be converted to a confidence map by using Gaussian blur.
The second section of the Affordance Sheet describes the datasets (RC1) used by the proposed solution, to detail their characteristics, share the link to the data, and the license informing about data permissions. We recommend future benchmarks to also release a detailed description on how to use and visualize data so that researchers can get acquainted with the format. Moreover, we recommend that future benchmarks evaluate models under different conditions, such as generalization to different object instances, object categories, object poses, backgrounds, and clutter. Benchmarks of models for tasks different from visual affordance prediction, such as COCO for object detection and instance segmentation [127] release only the training and validation sets while keeping a private testing set to not bias the designer of the architecture [127]. The availability of a testing set can lead researchers to make changes aimed at improving performance scores rather than formulating contributions that advance the field.
The third section highlights the model characteristics (RC2, RC3) integrating information in the model card (if available). Providing model cards [23], along with its implementation and trained weights, helps detail the description of models supporting other researchers to build upon. When not available, we encourage the re-implementation and retraining of the models as a contribution for the community (e.g. a previous work re-implemented, retrained, and released models for affordance detection and segmentation due to the lack of available models [4, 59]). As recommended for datasets, we encourage providing a link to the trained model’s weights and a license detailing the allowed uses. Without a license, the code is automatically protected by copyright, hence other researchers can not directly use the method implementation to reproduce results or as baseline, as for some previous works [56, 80, 86, 89, 92].
By providing the details of the experimental setup to train and evaluate methods (RC4), the fourth section of the Affordance Sheet is fundamental to correctly use previous methods and develop a solution under the same conditions. Setup conditions include pre-processing and post-processing information such as data splits, resize procedures, data normalisation, and hyper-parameters choice. The lack of these details can result in models with significantly different parameters, and hence leading to unfair comparisons with previous works.
The fifth section of the Affordance Sheet focuses on the performance measures (RC5), the criteria used to validate and compare methods with previous solutions. Providing a stand-alone toolkit implementing the performance measures ensures the replicability of the results across different works while including new methods. For visual affordance prediction, we recommend evaluating the performance of models using more than one measure to provide a more comprehensive analysis while identifying different aspects and limitations of the models. For example, in affordance segmentation, precision focuses on how many of the predicted pixels have the correct class and recall emphasizes how many of the annotated pixels are correctly predicted. Therefore, computing more than one score (and avoiding using a single score aggregating multiple performance measures) reduces the risk of drawing misleading conclusions that are based only on partial results.
The last section describes the validation of the method through a robotic setup. In previous works, few of the methods were validated using a robotic platform [6, 12, 77, 87, 108]. Unlike previous sections of the Affordance Sheet, the robot validation depends on the availability of a robot. When a robot experiment can be performed, we recommend reporting the characteristics of the setup, the robotic hand specifics, and the description of the experiment in terms of object and conditions. This transparent reporting allows researchers to assess methods using a common platform.
VI Future directions
In this section we discuss unexplored directions: estimating object physical properties, integration with AI agents, scaling datasets size, and benchmarking models performance.
Object physical properties. Relating affordance prediction and estimation of object physical properties is far from easy. Humans have different ways of grasping objects depending on the action they want to perform and how the object properties (e.g. mass) influences the action through the physics of the interaction [128]. Estimating object physical properties only from images might be too complex, and other modalities, such as language, audio, and haptic, could be included in our proposed formulation [129]. Multimodal models have shown better generalisation to novel and different object categories in tasks such as open-vocabulary object detection [130] and segmentation [131]. Language can be processed to select the most appropriate grasp for the task [132]. Audio could complement the visual modality when the appearance of the object is not reliable, e.g. an opaque container whose content is not visible [133]. Haptic could provide a feedback on the force that the agent applies on the object [134].
AI agents, human-in-the-loop, and VLA models. An AI agent [135] integrating visual affordance requires steps such as understanding (perception), reasoning (relating affordances, objects, and physical properties, conditioned to the task to accomplish), planning (actuation to accomplish the task), and recovering from errors. Learning to predict visual affordances for hand-object interactions can benefit from human demonstrations of the actions to perform, in the same way humans prompt models with examples showing how to solve tasks [136]. Our formulation can be extended to include the feedback from a person at different stages (human-in-the-loop) [137] to correct the prediction mistakes or also to inject task specific knowledge in the process. Another research direction is the conditioning of end-to-end models with affordance [138, 139]. These end-to-end methods do not explicitly model object affordances, and require thousands of demonstration data during training to generalise to different objects. Integrating affordance information coming from our formulation in end-to-end methods can improve spatial reasoning and generalisation to unseen tasks [140].